# version 0.6

### v3

@@ 2.3 Intractability of Equilibrium, Stabilizability, and Controllability
Figure 2.6: Julia set
Figure 2.7: Lorenz attractor
@@ 2.3.1 Computational Hardness of Stabilizability
We show that this problem can be reduced from the 3SAT problem, a known NP-hard problem. This implies that determining the stabilizability of a dynamical system is NP-hard in general.
@@
$x_{t+1}=x_{t}(1-\varepsilon+p_{F}(\mathbf{u}_{t}))+p_{F}(\mathbf{u}_{t}),$
with $p$ being the polynomial family we just constructed, and $\varepsilon$ being an arbitrarily small constant. For the system to reach the null state from an arbitrary start state, we would need to verify if there is an assignment of $\mathbf{u}$ such that $p_{F}(\mathbf{u})=0$, which we just proved is NP-hard. ∎
@@ 2.5 Exercises
2.1. Propose a computationally efficient representation for the problem of stabilizability. Prove that for your proposal, the problem of determining whether a given dynamical system is stabilizable belongs to the class NP. Thus, together with the material of this chapter, conclude that stabilizability is NP-complete.
@@ 3.1 Reinforcement Learning
Reinforcement learning is a sub-field of machine learning that formally models the setting of learning through interaction in a reactive environment. This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling. In reinforcement learning, we have an agent and an environment. The agent observes its position (or state) in the environment and takes actions that transition it to a new state. The environment looks at the agent’s state and hands out rewards based on a hidden set of criteria.
Figure 3.1: Source: Reinforcement Learning: An Introduction. Sutton & Barto, 2017.
@@
The reader can reason about situations where such an assumption is reasonable. For example, in the inverted pendulum setting described in section 2.1.3, the assumption clearly holds. It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onwards.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel & Norvig [RN02]. In this example, a robot starts in the bottom leftmost square, and its goal is to reach the $+1$ reward.
Figure 3.2: Source: [RN02].
In this example, the robot has controls that allow it to move left / right / forward / backward. The controls, however, are not reliable, and with $0.8$ probably the robot moves to wherever it intends. With the remaining $0.2$ probability the robot moves sideways, $0.1$ probability to each side.
@@
For example, if the rewards elsewhere are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
@@
On the other hand, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in the following.
@@ 3.2.1 Markov Chains
It consists of a set of states and a transition matrix describing the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. Figure 3.2.1 describes a Markov Chain corresponding to a weather process.
@@
$\mathbf{p}_{t+1}=\mathbf{p}_{t}P=\mathbf{p}_{0}P^{t},$
@@
A stationary distribution does not always exist nor is it always unique. Existence, uniqueness, and convergence to stationarity are the subject of Ergodic theory, a rich and well-studied branch of mathematics. For our discussion, we are interested in sufficient conditions for stationary distribution to exist. They are:
@@
    Aperiodicity. A Markov Chain is a-periodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
@@
A interesting special case of Markov chains is reversible Markov Chain.
@@
A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain that is reversible in the above definition.
@@ A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
When the discount factor that is strictly less than one, $G_{t}$ is well defined. In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that a stationary distribution for the underlying Markov chain exists, denoted by $\boldsymbol{\pi}$.
@@
The value function tells how valuable a state as measure by the sum of its future, discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@ A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
    $R$ is a reward function that gives each transition and and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$. We henceforth assume w.l.o.g. that $R\in[0,1]^{S\times A}$, where we define
@@
The fact that there exist an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
@@ 3.3 The Bellman Equation
The Bellman equation for any policy $\pi$ relates the value function for that policy to itself via one step of application for that policy. Mathematically, this amounts to the following
@@
    $P^{\pi}_{ss^{\prime}}=\sum_{a}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
@@ 3.4 The value iteration algorithm
The value iteration algorithm start from an arbitrary solution and improves it until a certain measure of optimality is reached. A formal description is given in Algorithm 2. The performance guarantee of Value Iteration, as well as its analysis, is particularly elegant and do not require expertise in mathematical optimization theory.
@@ 3.6 Exercises
3.1. In this exercise we give a guided proof of the ergodic theorem. Consider a given irreducible and a-periodic Markov chain with transition matrix $P$.
@@
C. Consider an abirtrary distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain $\mathbf{p}_{t}=P^{t}\mathbf{p}_{0}$. Prove that $\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
@@
3.4. Let Markov chain $P$ and distribution ${\boldsymbol{\pi}}$ be reversible for it, i.e.
@@
Our treatment in this chapter is restricted to linear dynamical systems (LDS): dynamical systems that evolves linearly according to the equation
@@
Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories.
A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time. These are called linear time-varying (LTV) systems. The class of LTV systems can be used to express much more general dynamics via linearization, as we show next.
4.1 General Dynamics as LTV Systems
@@
To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$ around some anchor point $(\bar{x}_{t},\bar{u}_{t})$:
@@
$\approx f(\bar{x}_{t},\bar{u}_{t})+\underbrace{\frac{\partial}{\partial x}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\triangleq A_{t}}(x_{t}-\bar{x}_{t})+\underbrace{\frac{\partial}{\partial u}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\triangleq B_{t}}(u_{t}-\bar{u}_{t})+w_{t}.$
With an appropriate redefinition of $w_{t}$, this simplifies to
$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t},$
where An analogous approximation holds for the higher-dimensional case. The Taylor expansion is
$f(\mathbf{x},\mathbf{u})\approx f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}}),$
@@ 4.2 Stabilizability of Linear Systems
although much of our discussion can be generalized to LTV systems. Exercise 4.6 asks the reader to prove this.
@@ 4.2.1 Efficient Verification of Stabilizability
From Theorem 4.3, we know that stabilizability of a linear system is equivalent to the existence of a matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)^{\top}P(A+BK)}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to exists of matrices ${K}$ and ${P}$ such that
@@ 4.4 Quantitative Definitions
In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control. These finite time bounds naturally depend on how difficult it is for a given system to stabilize. Thereforem we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
@@ 4.4.1 Stabilizability of Time-Invariant LDS
Consider a linear controller ${K}$ for the LTI system $({A,B})$. If there exists $\kappa>0$ and $\delta\in[0,1)$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
@@ 4.4.3 Controllability of time-invariant systems
Clearly, such a system is nominally controllable, as indeed is any one-dimensional system with non-zero $B$. However, to reach a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of size at least $1/\varepsilon$. Thus, such a system may be said to be barely controllable. This difficulty of controllability is captured in the definition below, which requires not only that $K_{k}(A,B)$ be full row-rank, but also that such a statement is robust to small changes in $A,B$.
@@
$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B))^{+}\|\leq\kappa.$
@@ 4.4. Prove that a stochastic linear dynamical system with Gaussian perturbations
B. For a time-invariant linear system, prove that there exists $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$. In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
C. Construct a time-varying linear systems that is BIBO stable, yet does not have the exponential decay properties required in Lemma 4.9.
@@
In the chapters thus far we have studied more basic notions than control, namely solution concepts of dynamical systems such as stabilizability and controllability. Even for these more basic problems, we have seen that they are computationally intractable in general.
@@
5.1 The Linear-Quadratic Regulator
The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs. Formally,

### v4

@@ 2.3 Intractability of Equilibrium, Stabilizability, and Controllability
Figure 2.6: Julia set. Souce: Wikipedia.
Figure 2.7: Lorenz attractor. Source: Wikipedia.
@@ 2.3.1 Computational Hardness of Stabilizability
We show that this problem can be reduced to the 3SAT problem, a well-known NP-hard problem. This implies that determining the stabilizability of a dynamical system is NP-hard in general.
@@
$x_{t+1}=(1-\varepsilon)x_{t}+p_{F}(\mathbf{u}_{t}),$
with $p$ being the polynomial family we just constructed and $\varepsilon>0$ an arbitrarily small constant. For the system to reach the null state from an arbitrary start state, we would need to verify if there is an assignment of $\mathbf{u}$ such that $p_{F}(\mathbf{u})=0$, which as we just proved is NP-hard. ∎
@@ 2.5 Exercises
2.2. Let $A\in\mathcal{R}^{d\times d}$ be a diagonalizable matrix. Consider the following dynamical system:
$\mathbf{x}_{t+1}=\frac{A\mathbf{x}_{t}}{\|\mathbf{x}_{t}\|}.$
What are the stable and attractive equilibrium points $\mathbf{x}\in\mathcal{R}^{d}$ of the system? Are they also globally attractive? Your answer may depend on $A$.
2.3. Let $\alpha,\beta,\gamma>0$. Does the SIR-model, as it is described in this chapter, define a stabilizable dynamical system (for an appropriate and fixed definition of the “zero state”)? What about controllability? We assume that only starting states with the property $S_{0}+I_{0}+R_{0}=1$ are allowed, and that $S_{t},I_{t},R_{t},u_{t}\geq 0$ has to hold at all times.
2.4. Consider the double integrator system:
$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+Bu_{t}+\mathbf{w}_{t},$
where
$A=\begin{bmatrix}1&\Delta\\ 0&1\end{bmatrix},\quad B=\begin{bmatrix}0\\ 1\end{bmatrix},\quad\mathbf{x}_{t}=\begin{bmatrix}x_{t}\\ v_{t}\end{bmatrix}\in\mathbb{R}^{2},\quad u_{t}\in\mathbb{R},\quad\Delta>0.$
The goal is to ensure the system reaches the desired state $[x^{*},0]^{\top}$ as $t\to\infty$ from any initial condition.
  1. 1.
    Prove that $(A,B)$ is controllable. That is, show that for any initial state and any target state, there exists a finite sequence of inputs that transitions the system from the initial state to the target state.
  2. 2.
    Design a stabilizing state-feedback law
[…]
    with $K=[k_{1}\;k_{2}]$, that places all eigenvalues of $A-BK$ strictly inside the unit circle. Find conditions on $\Delta,k_{1},k_{2}$ that ensure all eigenvalues of $A-BK$ are real and lie strictly between $-1$ and $1$, thus guaranteeing $\mathbf{x}_{t}\to[x^{*},0]^{\top}$.
@@ 3.1 Reinforcement Learning
Reinforcement learning is a sub-field of machine learning that formally models the setting of learning through interaction in a reactive environment. This is in contrast to learning from examples, as in supervised learning, or inference according to a probabilistic model, as in Bayesian modeling. In reinforcement learning, we have an agent and an environment. The agent observes its position (or state) in the environment and takes actions that transition it to a new state. The environment looks at the agent’s state and hands out rewards based on a set of criteria hidden from the agent.
Figure 3.1: Source: Wikipedia.
@@
The reader can imagine situations where such an assumption is reasonable. For example, in the inverted pendulum setting described in section 2.1.3, the assumption clearly holds. It is not significant for the control problem how a certain state, describing the angle and the angular velocity, was reached, in order to optimally control the pendulum from this state onward.
@@ 3.1.1 A Classical Example: Bubbo’s World
Before diving into the formalism of reinforcement learning, we start with a classical example from the text of Russel and Norvig [RN02]. In this example, a robot starts in the bottom leftmost square, and its goal is to reach the $+1$ reward. The blocks $+1,-1$ represent terminal states, meaning that the game ends when the robot reaches them. The block ‘X’ represents an obstacle through which the robot cannot travel.
StartX+1-1
Figure 3.2: A visual representation of Bubbo’s world.
In this example, the robot has controls that allow it to move left / right / forward / backward. The controls, however, are not reliable, and with $0.8$ probability the robot moves to wherever it intends. With the remaining $0.2$ probability the robot moves sideways, that is, with probability $0.1$ to each side.
@@
For example, if the rewards at all other states are slightly negative, Bubbo wants to reach the terminal state with a persistent $+1$ reward, as given in the following figure.
@@
On the other hand, if the rewards are very negative, Bubbo just wants to reach any terminal state as soon as possible, as in figure 3.4.
@@ 3.2.1 Markov Chains
It consists of a set of states and a transition matrix describing the probability of moving from one state to another. The probabilities of moving between states are the weights on the edges of the graph. The transition matrix is of dimension $S\times S$, where $S$ is the number of states, and in entry $i,j$ we have the probability of moving from state $i$ to state $j$. Figure 3.5 describes a Markov Chain corresponding to a weather process with a corresponding transition matrix $P$ as below
@@
Figure 3.5: A graphical description of a Markov decision process. Each state is the weather for a given day and the transition probabilities describe the chance of weather for the next day.
@@
$\mathbf{p}_{t+1}=\mathbf{p}_{t}P=\mathbf{p}_{0}P^{t+1},$
@@
A stationary distribution does not always exist nor is it always unique. Existence, uniqueness, and convergence to stationarity are the subject of Ergodic theory, a rich and well-studied branch of mathematics. For our discussion, we are interested in sufficient conditions for stationary distributions to exist. They are:
@@
    Aperiodicity. A Markov Chain is aperiodic if and only if $d(s_{i})=1$ for all states $s_{i}\in S$, where
@@
An interesting special case of Markov chains is the reversible Markov chain.
@@
A Markov chain is said to be reversible if there exists some distribution with respect to which the Markov chain is reversible.
@@ A Markov Reward Process is a tuple $(S,P,R,\gamma)$, where:
When the discount factor is strictly less than one, $G_{t}$ is well defined. In the case $\gamma=1$, an alternative definition of $G$ is presented below, assuming that a stationary distribution for the underlying Markov chain exists, denoted by $\boldsymbol{\pi}$.
@@
The value function tells how valuable a state is as measured by the sum of its future, discounted rewards. One way to compute the value function is by using the Bellman equation for the Markov Reward Process, which is the following recursive equation:
@@ A Markov Decision Process is a tuple $(S,P,R,A,\gamma)$ such that:
    $R$ is a reward function that gives each transition and action a reward $R_{s,s^{\prime},a}$ which is a reward for moving from $s$ to $s^{\prime}$ using action $a$. We henceforth assume w.l.o.g. that $R\in[0,1]^{S\times A}$, where we define
@@
The fact that there exists an optimal policy that is optimal regardless of the starting state in a MDP is a consequence of the Markovian nature of the transition dynamics.
@@ 3.3 The Bellman Equation
The Bellman equation for any policy $\pi$ relates the value function for that policy to itself via one step of application for that policy. Mathematically, this amounts to the following:
@@
    $P^{\pi}_{ss^{\prime}}=\sum_{a\in A}P^{a}_{ss^{\prime}}\cdot\pi(a|s).$
@@ 3.4 The Value Iteration Algorithm
The value iteration algorithm starts from an arbitrary solution and improves it until a certain measure of optimality is reached. A formal description is given in Algorithm 2. The performance guarantee of Value Iteration, as well as its analysis, is particularly elegant and does not require expertise in mathematical optimization theory.
@@ 3.6 Exercises
3.1. In this exercise we give a guided proof of the ergodic theorem. Consider a given irreducible and aperiodic Markov chain with transition matrix $P$.
@@
C. Consider an abirtrary distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain $\mathbf{p}_{t}=\mathbf{p}_{0}P^{t}$. Prove that $\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
@@
3.4. Let Markov chain $P$ be reversible with respect to stationary distribution ${\boldsymbol{\pi}}$, i.e.
@@
Our treatment in this chapter is restricted to linear dynamical systems (LDS): dynamical systems that evolve linearly according to the equation
@@
Figure 4.1: Four different deterministic linear time-invariant dynamical systems represented via their induced vector fields and a few chosen trajectories. Source: Wikipedia.
A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time. These are called linear time-varying (LTV) systems. The class of LTV systems can express more general dynamics via linearization, as we show next.
4.1 Approximating General Dynamics via LTV Systems
@@
To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$, which is twice differentiable, around an anchor point $(\bar{x}_{t},\bar{u}_{t})$:
@@
$=f(\bar{x}_{t},\bar{u}_{t})+\underbrace{\frac{\partial}{\partial x}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}A_{t}}(x_{t}-\bar{x}_{t})+\underbrace{\frac{\partial}{\partial u}f(x,u)\bigg{|}_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}B_{t}}(u_{t}-\bar{u}_{t})+w_{t},$
where $w_{t}$ is the approximation error, or remainder, which by the mean value theorem behaves as $w_{t}\sim(x_{t}-\bar{x}_{t})^{2}+(u-\bar{u}_{t})^{2}$. Thus, this simplifies to
$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t}.$
An analogous approximation holds for the higher-dimensional case. The Taylor expansion is
$f(\mathbf{x},\mathbf{u})=f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}})+w_{t},$
@@ 4.2 Stabilizability of Linear Systems
although much of our discussion can be generalized to LTV systems.
@@ 4.2.1 Efficient Verification of Stabilizability
Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and it holds that $\rho((A+BK)^{\top})=\rho(A+BL)$. Now using this fact and Theorem 4.3, we know that stabilizability of a linear system is equivalent to the existence of a matrices ${K}$ and ${P}$ such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to existence of matrices ${K}$ and ${P}$ such that
@@ 4.4 Quantitative Definitions
In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control. These finite time bounds naturally depend on how difficult it is for a given system to stabilize. Therefore, we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
@@ 4.4.1 Stabilizability of Time-Invariant LDS
Consider a linear controller ${K}$ for the LTI system $({A,B})$. If there exists $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
@@ 4.4.3 Controllability of time-invariant systems
Clearly, such a system is nominally controllable, as indeed is any one-dimensional system with non-zero $B$. However, to reach a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of size at least $1/\varepsilon$. Thus, such a system may be said to be barely controllable. This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ be full row-rank, but also that such a statement is robust to small changes in $A,B$.
@@
$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B))^{-1}\|\leq\kappa.$
In the definition above, the spectral norm of the inverse of $K(A,B)$ is bounded by $\kappa$. For example, in the simple one-dimensional system we presented above, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon}$, which indeed captures the complexity of controlling the system.
@@ 4.4. Prove that a stochastic linear dynamical system with Gaussian perturbations
B. For a time-invariant BIBO-stable linear system, prove that there exists $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$. In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
C. Construct a time-varying linear system that is BIBO stable, yet does not have the exponential decay properties required in Lemma 4.9.
@@
In the chapters thus far, we have studied more basic notions than control, namely solution concepts of dynamical systems such as stabilizability and controllability. Even for these more basic problems, we have seen that they are computationally intractable in general.
@@
5.1 The Linear Quadratic Regulator
The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs. For simplicity, we mostly consider the time-invariant version, where the linear dynamical system is fixed, and the costs are also fixed quadratic, as given formally below: